During 2017, Flint Corporation reported net sales of $2000000, net income of $860000, and depreciation expense of $150000. Flint also reported beginning total assets of $1010000, ending total assets of $1340000, plant assets of $590000, and accumulated depreciation of $750000. Flint’s asset turnover (rounded) is

a. 1.70 times.

b. 1.98 times

c. 0.78 times.

d. 1.49 times.
Answer #1

Assets turnover = Net Sales/Average assets

Average assets = (1010000+1340000/2) = 1175000

Assets turnover = 2000000/1175000 = 1.70

So answer is a) 1.70 Times